# Codenames as a Benchmark for Large Language Models

### v1

@@ I Introduction
To explore the potential of LLMs for language-based games, we have selected the board game Codenames as our benchmark [13]. Codenames is a popular team-based party game that requires cooperation, natural language understanding, and epistemic reasoning abilities to play effectively. Players work in asymmetric two-person teams, where one player (codemaster) provides a single word clue that attempts to link a variety of other words together. The second player (guesser) must then select from a board of several possible words, those which they feel are most associated with the provided clue. The full rules for playing Codenames are described in Section II, for any readers who are unfamiliar with this game.
@@ II-A Codenames Rules
The following four sub-sections describe the complete rules for Codenames. The format and wording used for these rule descriptions matches the input prompt provided to all LLM agents within our updated Codenames AI framework. For additional clarification, Figure 1 shows an example board setup from the codemaster’s perspective. The guesser is presented the same board but with the identities (i.e., colours) of the words hidden.
@@ II-A1 Overview
Codenames is a word-based game of language understanding and communication. Players are split into two teams (red and blue), with each team consisting of a Codemaster and Guesser.
@@ II-A2 Setup
At the start of the game, the board consists of 25 English words. The Codemaster on each team has access to a hidden map that tells them the identity of all of the words (Red, Blue, Civilian or Assassin). The Guessers on each team do not have access to this map, and so do not know the identity of any words. Players need to work as a team to select all their words in as few turns as possible, while minimising the number of incorrect guesses.
@@ II-A3 Turns
The clue number must be greater than or equal to zero. The Guesser then selects from the remaining words on he board, based on which word is most associated with the Codemaster’s clue. The identity of the selected word is then revealed to all players. If the Guesser selected a word that is their team’s colour, then they may get to select another word. The Guesser must always make at least one guess each turn, and can guess up to one word more than the number provided in the Codemaster’s clue. The only exception to this is if the Codemaster’s clue number is zero, then there is no limit on the maximum number of guesses. If a Guesser selects a word that is not their team’s colour, their turn ends. The Guesser can choose to stop selecting words (ending their turn) any time after the first guess.
@@ II-A4 Ending
Play proceeds, passing back and forth, until one of three outcomes is achieved:
  • All of the words of your team’s colour have been selected (your team wins).
  • All of the words of the other team’s colour have been selected (your team loses).
  • Your team selects the assassin word (your team loses).
@@ II-C Framework Limitations
Another rule from the original Codenames game relates to the penalty for invalid clues. In the official rulebook, if the codemaster gives an invalid clue their turn ends immediately and the other team’s codemaster gets to identify one of their team’s words for free. However, the imperfect nature of LLMs means that they have an increased risk of providing a clue in an invalid format (such as providing additional text in their response). We therefore decided to relax this rule, and instead simply ask the codemaster to try again if their original response was invalid. However, If the codemaster fails to provide a valid clue 10 times in a row, then a default empty string is chosen as the clue with the clue number being set to 1. Likewise, if the guesser fails to guess a valid word on the board 10 times in a row, then one of the remaining words is chosen at random. During our experiments, no codemaster ever failed to give a valid clue within this defined limit. However, the guesser agent would very occasionally fall into a repeated loop of providing a response that wasn’t on the board and would eventually have a word selected at random (although this happened only a handful of times across all experiments).
@@ III-B Reasoning Skills in Codenames
Lastly, Codenames also requires cooperative and epistemic reasoning. Both the codemaster and the guesser have to play with the each other in order to win, despite not being able to communicate outside of their defined actions. These Partial Information, Restricted Communication, Cooperative (PIRCC) games have recently garnered interest in the AI literature due to the complexity of creating capable AI players, and humans’ current superior performance to many AI [28]. To perform in these settings effectively, each agent needs to internalise what their teammate, and to a lesser extent the opposing team, may or may not think based on their actions (i.e., what is their current mental model). To perform better at the game, this reasoning needs to be applied when inducing suitable clues (codemaster) and deducing words on the board (guesser). This reasoning also allows players to adapt their strategy and playstyle based on the performance and actions of their teammate. For example, a player may give a clue about medieval history because their teammate is knowledgeable on the subject, or a highly skilled codemaster may need to reduce their clue complexity when paired with a novice guesser. LLMs may be able to use this type of reasoning to play Codenames better than other more traditional agents, having previously demonstrated some theory of mind capabilities [16, 29], although there remains debate over whether this reflects true possession of a theory of mind [24, 25].
@@ IV-A1 LLM Agents
However, it became apparent during preliminary testing that many of the smaller models were not able to correctly adhere to the game’s specified rules and output format, resulting in repeated invalid clues or guesses that frequently devolved into pure random play. As such, only the six LLMs highlighted in bold (o1-preview, o1-mini, GPT-4o, Gemini-1.5, Sonnet-3.5 and Llama-3.1) were able to consistently follow the game’s rules. Scores for the other LLMs were significantly worse because of this, and we thus chose to only present our findings and analysis for these six higher performing LLMs.
@@ IV-A2 Word-Vector Agents
One crucial limitation of these word-vector agents is that, due to the fact that their approaches utilise a pre-defined corpus of words for determining suitable clues / guesses, they are unable to interpret any words which are not present in their training set. Preliminary testing found that these agents will often produce errors when paired with LLM teammates, which have a much wider ranging vocabulary from which to select clues. Because of this, we were unable to produce reliable performance results for games with both LLM and Word-Vector agents, and have instead chosen to evaluate each agent group separately.
@@ IV-B2 Two Teams (competitive / cooperative)
IV-C Evaluation Procedure
@@
IV-C1 Single Team
This experiment compared the performance of different LLM and word-vector agent combinations, acting as both the codemaster and guesser, for the single team version of Codenames. Due to the previously mentioned issues with word-vector agents failing to understand clues provided by LLM agents, these two groups of agents were evaluated separately. In addition, due to the high usage costs currently required to run OpenAI’s recent o1 models, the o1-preview and o1-mini LLMs were only evaluated when paired with themselves.
IV-C2 Two Teams
@@
Regarding the costs associated with replicating these experiments, while the Llama models are open-source, the GPT, Gemini and Claude models are currently only available using a paid service offered by each provider. As of the time of writing, the costs incurred as a result of these experiments total approximately \$514 USD for GPT (OpenAI), \$141 USD for Gemini (Google Deepmind), and \$177 USD for Claude (Anthropic). Please note that this does not include the additional subscription fee required to access the OpenAI API. The bulk of these costs was attributed to the new o1-preview model provided by OpenAI, which cost \$288 USD alone for running 100 trials of the single team game version. While costs for using these models are expected to reduce over time, this can still be a prohibitively expensive requirement for any students or self-funded researchers wishing to experiment with these same models.
@@ V-A Single Team Version
  • Loss: Percentage of games that ended in a loss (i.e., selecting the assassin word).
@@
Mean Median Min Std Dev Loss Blue avg(stdev) Civillian avg(stdev) Clues avg(stdev) Guesses avg(stdev) Stop Early Stop Late
o1-preview - o1-preview 8.41 6 4 6.53 13% 0.84 (0.83) 1.20 (1.08) 1.95 (0.69) 1.88 (0.70) 11.0% 14.2%
o1-mini - o1-mini 11.10 7 4 7.93 24% 1.74 (1.19) 1.80 (1.24) 2.09 (0.70) 1.91 (0.81) 11.6% 16.0%
GPT-4o - GPT-4o 10.58 8 6 6.41 16% 0.96 (0.97) 1.25 (1.13) 1.51 (0.60) 1.45 (0.56) 0.0% 0.0%
GPT-4o - Gemini-1.5 9.81 8 6 5.21 10% 1.19 (1.0) 1.01 (0.88) 1.50 (0.60) 1.39 (0.56) 6.1% 0.5%
GPT-4o - Sonnet-3.5 10.50 8 6 6.20 15% 1.26 (1.16) 1.13 (0.86) 1.52 (0.61) 1.45 (0.57) 2.0% 1.6%
GPT-4o - Llama-3.1 11.17 9 5 6.43 17% 1.43 (1.25) 1.18 (0.94) 1.49 (0.60) 1.41 (0.56) 3.1% 3.8%
Gemini-1.5 - GPT-4o 10.95 8 5 6.55 17% 1.42 (1.23) 1.35 (1.14) 1.62 (0.72) 1.49 (0.64) 0.3% 0.0%
Gemini-1.5 - Gemini-1.5 10.38 8 5 6.08 14% 1.36 (1.06) 1.01 (0.88) 1.63 (0.73) 1.46 (0.62) 3.9% 0.2%
Gemini-1.5 - Sonnet-3.5 11.18 8 5 6.84 19% 1.42 (1.08) 1.15 (0.97) 1.62 (0.72) 1.51 (0.65) 2.9% 2.8%
Gemini-1.5 - Llama-3.1 11.28 9 5 6.46 17% 1.76 (1.45) 1.51 (1.10) 1.65 (0.71) 1.48 (0.63) 3.7% 1.8%
Sonnet-3.5 - GPT-4o 12.34 7 5 8.39 30% 1.94 (1.05) 1.87 (1.27) 2.04 (0.59) 1.88 (0.58) 3.7% 0.0%
Sonnet-3.5 - Gemini-1.5 10.55 8 5 6.89 18% 1.90 (1.24) 1.74 (1.06) 2.03 (0.61) 1.75 (0.64) 18.1% 0.0%
Sonnet-3.5 - Sonnet-3.5 11.01 8 5 7.55 22% 1.97 (1.15) 1.81 (1.00) 2.02 (0.60) 1.84 (0.61) 6.9% 0.6%
Sonnet-3.5 - Llama-3.1 11.31 8 5 7.20 21% 1.95 (1.34) 1.66 (1.13) 1.99 (0.61) 1.66 (0.65) 21.3% 0.5%
Llama-3.1 - GPT-4o 10.56 9 6 5.51 12% 1.13 (0.99) 0.95 (0.86) 1.34 (0.53) 1.29 (0.50) 0.3% 0%
Llama-3.1 - Gemini-1.5 10.54 9 6 5.29 11% 1.01 (0.94) 1.04 (0.93) 1.32 (0.51) 1.29 (0.48) 0.8% 0.3%
Llama-3.1 - Sonnet-3.5 10.79 9 6 5.66 13% 1.10 (0.95) 1.01 (0.99) 1.34 (0.53) 1.30 (0.50) 0.6% 1.6%
Llama-3.1 - Llama-3.1 10.18 9 6 4.66 8% 1.28 (1.10) 1.13 (1.02) 1.30 (0.51) 1.30 (0.50) 0.1% 5.7%
Word2Vec - Word2Vec 6.81 6 4 0.97 0% 0.0 (0.0) 0.0 (0.0) 1.46 (0.64) 1.46 (0.64) 0.0% 0.0%
Word2Vec - GloVe 12.35 9 5 7.14 22% 1.68 (1.78) 1.38 (1.53) 1.53 (0.65) 1.38 (0.58) 0.0% 0.0%
Word2Vec - Combined 10.64 8 5 6.72 16% 0.99 (1.48) 1.05 (1.30) 1.48 (0.64) 1.39 (0.60) 0.0% 0.0%
GloVe - Word2Vec 12.21 9 5 7.52 24% 2.03 (1.53) 1.45 (1.19) 1.74 (0.76) 1.54 (0.69) 0.0% 0.0%
GloVe - GloVe 5.24 5 4 0.79 0% 0.0 (0.0) 0.0 (0.0) 1.72 (0.75) 1.72 (0.75) 0.0% 0.0%
GloVe - Combined 5.42 5 4 0.89 0% 0.15 (0.39) 0.09 (0.29) 1.71 (0.75) 1.70 (0.75) 0.0% 0.0%
Combined - Word2Vec 11.18 8 5 7.37 21% 1.47 (1.30) 1.04 (1.07) 1.66 (0.74) 1.54 (0.69) 0.0% 0.0%
Combined - GloVe 5.86 6 4 2.18 1% 0.09 (0.32) 0.12 (0.36) 1.64 (0.73) 1.63 (0.72) 0.0% 0.0%
Combined - Combined 5.53 6 4 0.83 0% 0.0 (0.0) 0.0 (0.0) 1.63 (0.73) 1.63 (0.73) 0.0% 0.0%

### v2

@@ I Introduction
To explore the potential of LLMs for language-based games, we have selected the board game Codenames (Czech Games Edition, 2015) as our benchmark [13]. Codenames is a popular team-based party game that requires cooperation, natural language understanding, and epistemic reasoning abilities to play effectively. Players work in asymmetric two-person teams, where one player (codemaster) provides a single word clue that attempts to link a variety of other words together. The second player (guesser) must then select from a board of several possible words, those which they feel are most associated with the provided clue. The full rules for playing Codenames are described in Section II, for any readers who are unfamiliar with this game.
@@ II-A Codenames Rules
The following four sub-sections describe the complete rules for Codenames. For additional clarification, Figure 1 shows an example board setup from the codemaster’s perspective. The guesser is presented the same board but with the identities (i.e., colours) of the words hidden.
@@ II-A1 Overview
Codenames is a word-based game of language understanding and communication. Players are split into two teams (red and blue), with each team consisting of a Codemaster and Guesser. The red team always goes first.
@@ II-A2 Setup
At the start of the game, the board consists of 25 English words. The Codemaster on each team has access to a hidden map that tells them the identity of all of the words (Red, Blue, Civilian or Assassin). A standard map in Codenames has 9 red words, 8 blue words, 7 civilian words and 1 assassin word. The Guessers on each team do not have access to this map, and so do not know the identity of any words. Players need to work as a team to select all their words in as few turns as possible, while minimising the number of incorrect guesses.
@@ II-A3 Turns
The clue number must be greater than or equal to zero. The Guesser then selects from the remaining words on the board, based on which word is most associated with the Codemaster’s clue. The identity of the selected word is then revealed to all players. If the Guesser selected a word that is their team’s colour, then they may get to select another word. The Guesser must always make at least one guess each turn, and can guess up to one word more than the number provided in the Codemaster’s clue. The only exception to this is if the Codemaster’s clue number is zero, then there is no limit on the maximum number of guesses. If a Guesser selects a word that is not their team’s colour, their turn ends. The Guesser can choose to stop selecting words (ending their turn) any time after the first guess.
@@ II-A4 Ending
Play proceeds, passing back and forth, until one of two outcomes is achieved:
  • All of the words of a team’s colour have been selected (this team wins).
  • A team’s guesser selects the assassin word (this team loses).
SINK POOL KNIFE ALPS CONTRACT
CAR PLATE TRUNK WORM RULER
BELT CHINA PARACHUTE FIGHTER SPELL
PRESS LION JAM MAPLE BEACH
CHICK MOUTH EMBASSY LEMON SCHOOL
@@ II-C Framework Limitations
Another rule from the original Codenames game relates to the penalty for invalid clues. In the official rulebook, if the codemaster gives an invalid clue their turn ends immediately and the other team’s codemaster gets to identify one of their team’s words for free. However, the imperfect nature of LLMs means that they have an increased risk of providing a clue in an invalid format (such as providing additional text in their response). We therefore decided to relax this rule, and instead simply ask the codemaster to try again if their original response was invalid. However, If the codemaster fails to provide a valid clue 10 times in a row, then a default empty string is chosen as the clue with the clue number being set to 1. Likewise, if the guesser fails to guess a valid word on the board 10 times in a row, then one of the remaining words is chosen at random. During our experiments, no codemaster ever failed to give a valid clue within this defined limit. However, the guesser agent would very occasionally fall into a repeated loop of providing a response that wasn’t on the board and would eventually have a word selected at random (although this happened only a handful of times across all experiments). While this modification provides a more lenient version of the game’s rules, it is important to highlight that the creation of LLMs that can accurately adhere to their provided prompts is a strict requirement of reliable AI. Future versions of this benchmark may choose to remove this modification in order to emphasise this rule following aspect of LLM evaluation.
@@ III-B Reasoning Skills in Codenames
Lastly, Codenames also requires cooperative and epistemic reasoning. Both the codemaster and the guesser have to play with each other in order to win, despite not being able to communicate outside of their defined actions. These Partial Information, Restricted Communication, Cooperative (PIRCC) games have recently garnered interest in the AI literature due to the complexity of creating capable AI players, and humans’ current superior performance to many AI [28]. To perform in these settings effectively, each agent needs to internalise how their teammate, and to a lesser extent the opposing team, may interpret their actions (i.e., what is their current mental model). To perform better at the game, this reasoning needs to be applied when inducing suitable clues (codemaster) and deducing words on the board (guesser). This reasoning also allows players to adapt their strategy and playstyle based on the performance and actions of their teammate. For example, a player may give a clue about medieval history because their teammate is knowledgeable on the subject, or a highly skilled codemaster may need to reduce their clue complexity when paired with a novice guesser. LLMs may be able to use this type of reasoning to play Codenames better than other more traditional agents, having previously demonstrated some theory of mind capabilities [16, 29], although there remains debate over whether this reflects true possession of a theory of mind [24, 25].
@@ IV-A1 LLM Agents
However, it became apparent during preliminary testing that many of the smaller models were not able to correctly adhere to the game’s specified rules and output format, resulting in repeated invalid clues or guesses that frequently devolved into pure random play. As such, only the nine LLMs highlighted in bold (o1-preview, o1-mini, o3-mini, GPT-4o, Gemini-1.5, Sonnet-3.5, DeepSeek-R1, DeepSeek-V3 and Llama-3.1) were able to consistently follow the game’s rules. Scores for the other LLMs were significantly worse because of this, and we thus chose to only present our findings and analysis for these nine higher performing LLMs.
@@ IV-A2 Word-Vector Agents
One crucial limitation of these word-vector agents is that, due to the fact that their approaches utilise a pre-defined corpus of words for determining suitable clues / guesses, they are unable to interpret any words which are not present in their training set. Preliminary testing found that these agents will often produce errors when paired with LLM teammates, which have a much wider ranging vocabulary from which to select clues. Because of this, we were unable to produce reliable performance results for games with both LLM and Word-Vector agents on the same team, and have instead chosen to evaluate each agent group separately.
@@ IV-B2 Two Teams (competitive / cooperative)
IV-C LLM Agent Prompts
When an LLM agent is initialised within the Codenames AI framework it is first provided with an input prompt describing the game’s rules, using the same format and wording shown in section II.A, along with its team’s colour and whether it is playing as the codemaster or guesser (e.g., “You are playing the game Codenames as the Red Guesser”). When required to give a response during the game, each agent is prompted as follows:
IV-C1 Codemaster
The only situation where the codemaster needs to provide a response is when giving a clue (word and number) based on the current board state. The codemaster is provided with the remaining words on the board ordered by their identity (red, blue, civilian and assassin) along with an instruction to “provide a single word clue and number for the guesser in the following format ('pebble',2)”. The codemaster is also reminded that “the clue cannot be derived from or derive one of the words on the board”, and informed that they are to “stick to this format exactly and provide no additional text”. If the provided clue format is invalid or violates the “no derived words” rule, the codemaster is informed of this and asked to respond again.
IV-C2 Guesser
The guesser has two possible situations where they are required to give a response: selecting a word on the board that matches the provided clue, and deciding whether to continue guessing after a correct guess has been made. When required to select a word for their guess, the guesser is provided with all the remaining words on the board as well as the codemaster’s clue (e.g., “(pebble,2)”). The guesser is instructed to “select one of the remaining words that is most associated with this clue”, and informed that they “must select one of the remaining words and provide no additional text”. When deciding whether to continue guessing, the guesser is once again informed of the remaining words on the board and the codemaster’s clue. The guesser is also told how many words it has already picked this turn, and is then asked “Would you like to keep guessing? Answer only 'yes'or 'no'”. If an invliad response is provided in either of these cases, the guesser is informed of this and asked to respond again.
IV-D Evaluation Procedure
@@
IV-D1 Single Team
This experiment compared the performance of different LLM and word-vector agent combinations, acting as both the codemaster and guesser, for the single team version of Codenames. Due to the previously mentioned issues with word-vector agents failing to understand clues provided by LLM agents, these two groups of agents were evaluated separately. In addition, due to the high usage costs currently required to run OpenAI’s recent o1/o3 models, the o1-preview, o1-mini and o3-mini LLMs were only evaluated when paired with themselves. Likewise, the DeepSeek models were also only evaluated when paired with themselves, although this was due to imposed usage limits at the time of writing.
IV-D2 Two Teams
@@
Regarding the costs associated with replicating these experiments, while the Llama models are open-source, the GPT, Gemini and Claude models are currently only available using a paid service offered by each provider. As of the time of writing, the costs incurred as a result of these experiments total approximately \$538 USD for GPT (OpenAI), \$141 USD for Gemini (Google Deepmind), \$177 USD for Claude (Anthropic) and \$8.33 USD for DeepSeek. Please note that this does not include the additional subscription fee required to access the OpenAI API. The bulk of these costs was attributed to the new o1-preview model provided by OpenAI, which cost \$288 USD alone for running 100 trials of the single team game version. While costs for using these models are expected to reduce over time, this can still be a prohibitively expensive requirement for any students or self-funded researchers wishing to experiment with these same models.
@@ V-A Single Team Version
  • Loss: Percentage of games that ended in a loss. Note, for our single team results all losses were caused by selecting the assassin word (rather than selecting all of the blue words).
  • Mean (without loss): Mean score across all games that did not end in a loss (i.e., games where all red words were identified).
@@
Mean Median Min Std Dev Loss Mean (without loss) Blue avg(stdev) Civilian avg(stdev) Clues avg(stdev) Guesses avg(stdev) Stop Early Stop Late
o1-preview - o1-preview 8.41 6 4 6.53 13% 5.93 0.84 (0.83) 1.20 (1.08) 1.95 (0.69) 1.88 (0.70) 11.0% 14.2%
o1-mini - o1-mini 11.10 7 4 7.93 24% 6.71 1.74 (1.19) 1.80 (1.24) 2.09 (0.70) 1.91 (0.81) 11.6% 16.0%
o3-mini - o3-mini 9.70 6 4 7.60 19% 6.11 1.68 (1.05) 1.70 (1.23) 3.40 (1.69) 2.04 (0.95) 17.4% 0.4%
DeepSeek-R1 - DeepSeek-R1 11.24 7 4 8.11 25% 6.65 1.76 (1.28) 2.04 (1.18) 2.95 (1.09) 1.97 (0.85) 16.0% 0.0%
DeepSeek-V3 - DeepSeek-V3 10.59 8 5 6.42 16% 7.85 1.43 (1.09) 1.23 (1.04) 1.77 (0.72) 1.50 (0.52) 14.7% 0.0%
GPT-4o - GPT-4o 10.58 8 6 6.41 16% 7.83 0.96 (0.97) 1.25 (1.13) 1.51 (0.60) 1.45 (0.56) 0.0% 0.0%
GPT-4o - Gemini-1.5 9.81 8 6 5.21 10% 8.12 1.19 (1.0) 1.01 (0.88) 1.50 (0.60) 1.39 (0.56) 6.1% 0.5%
GPT-4o - Sonnet-3.5 10.50 8 6 6.20 15% 7.94 1.26 (1.16) 1.13 (0.86) 1.52 (0.61) 1.45 (0.57) 2.0% 1.6%
GPT-4o - Llama-3.1 11.17 9 5 6.43 17% 8.34 1.43 (1.25) 1.18 (0.94) 1.49 (0.60) 1.41 (0.56) 3.1% 3.8%
Gemini-1.5 - GPT-4o 10.95 8 5 6.55 17% 8.07 1.42 (1.23) 1.35 (1.14) 1.62 (0.72) 1.49 (0.64) 0.3% 0.0%
Gemini-1.5 - Gemini-1.5 10.38 8 5 6.08 14% 8.00 1.36 (1.06) 1.01 (0.88) 1.63 (0.73) 1.46 (0.62) 3.9% 0.2%
Gemini-1.5 - Sonnet-3.5 11.18 8 5 6.84 19% 7.94 1.42 (1.08) 1.15 (0.97) 1.62 (0.72) 1.51 (0.65) 2.9% 2.8%
Gemini-1.5 - Llama-3.1 11.28 9 5 6.46 17% 8.47 1.76 (1.45) 1.51 (1.10) 1.65 (0.71) 1.48 (0.63) 3.7% 1.8%
Sonnet-3.5 - GPT-4o 12.34 7 5 8.39 30% 6.91 1.94 (1.05) 1.87 (1.27) 2.04 (0.59) 1.88 (0.58) 3.7% 0.0%
Sonnet-3.5 - Gemini-1.5 10.55 8 5 6.89 18% 7.38 1.90 (1.24) 1.74 (1.06) 2.03 (0.61) 1.75 (0.64) 18.1% 0.0%
Sonnet-3.5 - Sonnet-3.5 11.01 8 5 7.55 22% 7.06 1.97 (1.15) 1.81 (1.00) 2.02 (0.60) 1.84 (0.61) 6.9% 0.6%
Sonnet-3.5 - Llama-3.1 11.31 8 5 7.20 21% 7.67 1.95 (1.34) 1.66 (1.13) 1.99 (0.61) 1.66 (0.65) 21.3% 0.5%
Llama-3.1 - GPT-4o 10.56 9 6 5.51 12% 8.59 1.13 (0.99) 0.95 (0.86) 1.34 (0.53) 1.29 (0.50) 0.3% 0%
Llama-3.1 - Gemini-1.5 10.54 9 6 5.29 11% 8.75 1.01 (0.94) 1.04 (0.93) 1.32 (0.51) 1.29 (0.48) 0.8% 0.3%
Llama-3.1 - Sonnet-3.5 10.79 9 6 5.66 13% 8.67 1.10 (0.95) 1.01 (0.99) 1.34 (0.53) 1.30 (0.50) 0.6% 1.6%
Llama-3.1 - Llama-3.1 10.18 9 6 4.66 8% 8.89 1.28 (1.10) 1.13 (1.02) 1.30 (0.51) 1.30 (0.50) 0.1% 5.7%
Word2Vec - Word2Vec 6.81 6 4 0.97 0% 6.81 0.0 (0.0) 0.0 (0.0) 1.46 (0.64) 1.46 (0.64) 0.0% 0.0%
Word2Vec - GloVe 12.35 9 5 7.14 22% 8.78 1.68 (1.78) 1.38 (1.53) 1.53 (0.65) 1.38 (0.58) 0.0% 0.0%
Word2Vec - Combined 10.64 8 5 6.72 16% 7.90 0.99 (1.48) 1.05 (1.30) 1.48 (0.64) 1.39 (0.60) 0.0% 0.0%
GloVe - Word2Vec 12.21 9 5 7.52 24% 8.17 2.03 (1.53) 1.45 (1.19) 1.74 (0.76) 1.54 (0.69) 0.0% 0.0%
GloVe - GloVe 5.24 5 4 0.79 0% 5.24 0.0 (0.0) 0.0 (0.0) 1.72 (0.75) 1.72 (0.75) 0.0% 0.0%
GloVe - Combined 5.42 5 4 0.89 0% 5.42 0.15 (0.39) 0.09 (0.29) 1.71 (0.75) 1.70 (0.75) 0.0% 0.0%
Combined - Word2Vec 11.18 8 5 7.37 21% 7.51 1.47 (1.30) 1.04 (1.07) 1.66 (0.74) 1.54 (0.69) 0.0% 0.0%
Combined - GloVe 5.86 6 4 2.18 1% 5.67 0.09 (0.32) 0.12 (0.36) 1.64 (0.73) 1.63 (0.72) 0.0% 0.0%
Combined - Combined 5.53 6 4 0.83 0% 5.53 0.0 (0.0) 0.0 (0.0) 1.63 (0.73) 1.63 (0.73) 0.0% 0.0%